HUBBLE SPACE TELESCOPE DAILY REPORT #2240 PERIOD COVERED: 0000Z (UTC) 10/29/98 - 0000Z (UTC) 10/30/98 Daily Status Report as of 303/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.3 Completed Four Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.4 Completed WF/PC-2 7497 (WF/PC-2 Observations of Astrophysically Important Visual Binaries) The WF/PC-2 was used to make observations of the visual binary variable star system Procyon. This system is a famous metal-deficient G dwarf for which accurate masses will lead to the star's helium content with cosmological implications. The observations were completed as planned, and no anomalies were reported. 1.5 Completed Three Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation II) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-1351 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.9 Completed WF/PC-2 6770 (Resolving the Nature of Nova Shells) The WF/PC-2 was used to make images of the post nova remnant HR-Del in order to study its surrounding shells structure. Past data has indicate a strong correlation between nova speed class and the ellipticity of the resulting remnants. The observations were completed as scheduled, and no problems were noted. 1.10 Completed STIS (CCD/MA1/MA2) 7346 (The Nature of the Bright 'Supersodt X-ray Source' RX J 0019+21) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to obtain decisive information on the true nature of RX J 0019+21, a bright "supersoft x-ray source". Two sets of echelle spectra at the quadrature phases of the binary orbit in order to determine the UV spectral energy distribution, to separate stellar and interstellar components, to search for information on the absorption of stellar or disk origin, to measure LineHell1160 emission lines, and to search for other emission/absorption lines. The proposal completed with no reported problems. 1.11 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As detailed in HSTAR 6746 and 2.1, the acquisition for the final iteration of this proposal failed at 302/192715Z after it had lost lock several times. The take data flag remained down, affecting all observations in this iteration. Otherwise, the observations were completed as planned, and no further problems were reported. 1.12 Completed STIS/CCD 8071 (HDF-South Flanking Fields) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the distant object SAO255267 as part of the HDF-South flanking field study. This program is designed to provide coverage around the HDF-South field and to provide a deep optical images of the NICMOS HDF-South field. The images were completed as planned, and no problems were reported. 1.13 Completed Three Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.14 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6837 (The Ultraviolet Properties of "Dwarf" Seyfert 1 Nuclei) The WF/PC-2 was used to make ultraviolet observations of the "dwarf", type one, Seyfert galaxy NGC-1161. This class of "dwarf" Seyfert populate the faint end of the local AGN luminosity function and hold important clues to the evolution of the entire AGN population. The observations were completed as scheduled, and no problems were reported. 1.16 Completed NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make observations of CALAR3. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. As detailed in HSTAR 6746 and 2.1, the acquisition for this proposal failed at 302/192715Z after it had lost lock several times. The take data flag remained down, affecting all observations. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 8 HSTAR 6746 describes a guide star acquisition failure occurring at 302/192715Z after several losses of lock. The proposals described in 1.11 and 1.16 were affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared five times. The STIS MCE-2 reset at 302/173700Z while in an SAA contour and while the high voltage was off. HSTAR 6745 was written. ROP NS-12 was executed to reset the STIS software error counter at 302/1825Z. The STIS MAMA-2 will be recovered at the time of the next high voltage turn on, 305/0747Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The "burn-in" of new PDB 53.C will begin today and continue throughout the weekend. Installation of this data base is expected to occur on Monday. Continuation of normal science observations and calibrations. /DMH